OpenStudy (anonymous):

Prove if gcd(a, b) = 1, then gcd(2a + b, a + 2b) is 1 or 3.

jhonyy9 (jhonyy9):

gcd(a,b)=1 so gcd(2a+b,a+2b)=gcd(3a,3b)=gcd(3(a,b))=1 or 3